OpenStudy (anonymous):

True or False. The coordinates of the midpoint of a segment on the coordinate plane are the products of the coordinates of the endpoints of that segment?

OpenStudy (lgbasallote):

Coordinate of midpoint: \(\huge (\frac{X_{1} + X_{2}}{2}, \frac{Y_{1} + Y_{2}}{2})\)

OpenStudy (anonymous):

that's a true statement :)

OpenStudy (anonymous):

@brandoduh: why do you think that is true? Please enlighten us :)

OpenStudy (anonymous):

the question is asking if you multiply the end points to get the midpoints, thats not what @lgbasallote is stating

OpenStudy (anonymous):

because you add x1 and x2 then you add y1 and y2. after that, you divide them by two. this is the midpoint formula

OpenStudy (lgbasallote):

so then..what is the answer to the question?

OpenStudy (anonymous):

its true :) thank you all

OpenStudy (anonymous):

It's not true. Lgba answer is right.

OpenStudy (lgbasallote):

i think either you misinterpreted the formula...or the question...the question is asking \(\LARGE (X_{1}) \times (X_{2}) = X_{M}\) where \(X_{M}\) is the midpoint of x \(\LARGE (Y_{1}) \times (Y_{2}) = Y_{M}\) where \(Y_{M}\) is the midpoint of y now do you see that in the midpoint formula?